# Approaching the Quantum Speed Limit in Quantum Gates with Geometric Control

### v1

@@ II Saturation of the MT-QSL bound for specific families of unitary dynamics
To illustrate the constraints imposed by the Mandelstam-Tamm QSL for unitary operators, we examine the potential saturation of this quantum speed limit in low-dimensional systems and discuss the brachistochrone approach within the QSL framework.
@@ II.1 Saturation of the MT-QSL in a 2D system
In a two-dimensional system, up to an arbitrary phase, a generic quantum state is fully defined by two angles, in agreement with the conventional Bloch sphere representation. As a result, when defining a general unitary operator, the mapping of the first state of the canonical basis corresponds to two degrees of freedom. The second basis state is necessarily mapped to an orthogonal state, but its phase can still be chosen independently. Therefore, quantum gates in a two-dimensional Hilbert space are parametrized by three degrees of freedom.
First, we examine the saturation of the QSL within a family of unitary dynamics parameterized by two control parameters $U(\theta(t),\varphi(t))=e^{-\frac{i}{2}\varphi(t)}|\psi_{+}(\theta(t))\rangle\langle 0|+e^{\frac{i}{2}\varphi(t)}|\psi_{-}(\theta(t))\rangle\langle 1|$ with the two orthogonal states $|\psi_{+}(\theta)\rangle=\cos(\theta/2)|0\rangle+\sin(\theta/2)|1\rangle$ and $|\psi_{-}(\theta)\rangle=-\sin(\theta/2)|0\rangle+\cos(\theta/2)|1\rangle$. This corresponds to a subset of the possible unitary gates in a two-dimensional system, but the extension to the more general case, which includes all possible unitary operators, is straightforward. Here, we aim to determine the dynamics of the parameters $(\theta(t),\varphi(t))$ such that the QSL is saturated at all time. The Hamiltonian is determined by $H(t)=i\hbar U^{\dagger}(t)\dot{U}(t)$ and is given by $H(t)=\frac{\hbar}{2}\left(\dot{\theta}\sigma_{y}+\dot{\varphi}\sigma_{z}\right)$ where $\sigma_{x,y,z}$ are the usual Pauli operators. We obtain $\Delta H=\frac{\hbar}{2}\sqrt{\dot{\theta^{2}}+\dot{\varphi}^{2}}.$ On the other hand, the overlap with the initial evolution operator is given by $\langle U_{0},U(t)\rangle=\cos\left(\frac{1}{2}\theta(t)\right)\cos\left(\frac{1}{2}\varphi(t)\right)$. By setting $\cos\phi=\langle U_{0},U(t)\rangle$, we find
@@
with $x=\cos\left(\frac{\theta(t)}{2}\right)$ and $y(t)=\cos\left(\frac{\varphi(t)}{2}\right).$ As $|y|\leq 1$, one finds $\sqrt{1-x^{2}}y=\sqrt{y^{2}-x^{2}y}\leq\sqrt{1-x^{2}y^{2}}$ and similarly $x\sqrt{1-y^{2}}\leq\sqrt{1-x^{2}y^{2}}$, which yields:
$\dot{\phi}\leq\frac{\dot{\theta}+\dot{\varphi}}{2}$ (18)
One has $\dot{\theta}+\dot{\varphi}\leq\sqrt{\dot{\theta}^{2}+\dot{\varphi}^{2}}$ by concavity of the function $x\rightarrow\sqrt{x^{2}+y^{2}}$ for any $y$, so that our result is consistent with the MT-QSL (9). Saturation occurs only if $x=0$ or $y=0$ at all times, meaning that either $\dot{\theta}=0$ at all times, or $\dot{\varphi}=0$ at all times. The former case corresponds to a trivial scenario where the unitary operator remains the identity throughout. Therefore, we must have $\dot{\varphi}=0$ at all times. We recover here the fact that the geodesics on a sphere coincide with its great arcs. Note that one may alternatively define an optimal QSL with respect to some unitary operator $U_{1}$ reached after a certain evolution. This example shows that saturating the QSL severely constrains the choice of possible quantum paths. To achieve QSL saturation, only one of the two angles $(\theta(t),\varphi(t))$ can evolve at a time.
@@ II.2 Investigation of the QSL limit in 3-level systems with Rodrigues formula
In Ref. [34], a method was proposed to control the dynamics of a three-level system by separating the evolution into two components: population changes, parameterized using Rodrigues’ rotation formula, and phase changes. This approach was used to inversely design the Hamiltonian of the three-level system, enabling the desired transition while ensuring allowed couplings and suppressing forbidden ones.
Similarly, we focus here on a three-state system and the family of unitary operators ${U_{r}(t)}$ that govern population transfer. Specifically, we consider mappings between quantum states with positive real-valued coefficients, noting $\mathcal{S}=\{|\psi\rangle|\forall n\in\{1,2,3\}\>\langle n|\psi\rangle\in\mathbf{R}^{+}\}$ the corresponding set of quantum states. In other words, if $|\psi(0)\rangle=a_{1}(0)|1\rangle+a_{2}(0)|2\rangle+a_{3}(0)|3\rangle$, where $a_{n}(0)\in\mathbb{R}^{+}$ for $n\in\{1,2,3\}$, then at any time $|\psi(t)\rangle=U_{r}(t)|\psi(0)\rangle$, the condition $\langle n|\psi(t)\rangle\in\mathbb{R}^{+}$ holds. Owing to their positivity, the unitary operators $U_{r}(t)$ correspond to rotations in a three-dimensional Euclidean space [34]. A 3D vector rotation is described by the following transformation:
$\mathbf{v}^{\prime}=\mathbf{v}\cos\theta+\mathbf{u}(\mathbf{u}\cdot\mathbf{v})(1-\cos\theta)+(\mathbf{u}\times\mathbf{v})\sin\theta\,$ (19)
where $\mathbf{v}$ is an arbitrary vector to which the transformation is applied, $\mathbf{v}^{\prime}$ its transformed image, $\mathbf{u}$ the rotation axis and $\theta$ the angle of rotation. The corresponding relation for the quantum evolution operator is given by
@@
where we have introduced $|u\rangle=\sum_{m}u_{m}|m\rangle$ $u$ is a real-valued vector of positive coefficient and unit norm - and $V=\sum_{i,j,k}\epsilon_{ijk}u_{i}|k\rangle\langle j|$, with $\epsilon_{ijk}$ the Levi-Civita symbol. Any evolution operator $U_{r}(t)$ that leaves invariant the subset $\mathcal{S}$, i.e. such that $U_{r}(t)(\mathcal{S})\subset\mathcal{S}$, can be expressed in this form.
Using this formalism, we investigate the saturation of the QSL. One of the conditions (7) indicates that the three operators ${U_{0},U(t),H(t)U(t)=i\dot{U}(t)}$ must lie within a common plane. Expressing $H(t)U(t)$ as $H(t)U(t)=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U(t)$ with $(\lambda_{1}(t),\lambda_{2}(t))\in\mathbb{C}^{2}$, we derive an alternative expression for the Hamiltonian: $H(t)=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U^{\dagger}(t)$. The Hermiticity condition of the Hamiltonian then requires that $U^{\dagger}(t)+U(t)=\lambda(t)\mathbb{I}$ for some complex number $\lambda(t)$. Since $V^{\dagger}=-V$, from Eq. (20) this condition simplifies to $\cos\theta\mathbb{I}+(1-\cos\theta)|u\rangle\langle u|=\lambda_{0}(t)\mathbb{I}$. Clearly, this is only satisfied if $\theta=0$, implying that no motion occurs.
@@ II.3 Quantum Brachistochrone for unitary evolution and the MT-QSL limit
As a matter of fact, this condition simply corresponds to a time-independent Schrödinger equation and is less stringent than the plane condition that arises for the MT-QSL bound, which imposes that $\dot{U}=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U(t)$.
Indeed, it is worth noticing that several examples of time-optimal unitary driving in a four-state systems, derived within the brachistochrone framework, are reported in Ref. [26] do not saturate the MT-QSL bound.
In the following example, we consider two interacting spins $1/2$, such that the system Hamiltonian takes the form $H(t)=-\sum_{j=x,y,z}J_{j}(t)\sigma_{j}^{1}\sigma_{j}^{2}+\sum_{a=1,2}B_{a}(t)\sigma_{z}^{a}$ with $\sigma^{1}_{j}:=\sigma_{1}\otimes\mathbf{1}_{2},$ $\sigma^{2}_{j}:=\mathbf{1}_{2}\otimes\sigma_{j}$ where $\sigma_{j}$ are the usual Pauli matrices for $j=x,y,z$, and the parameters $J_{j}(t),B_{a}(t)$ are free parameters determined by the brachistochrone equation. The time-optimal realization of a SWAP gate
@@
we obtain $|\langle\mathbb{I},U_{\rm QFT}\rangle|=1/\sqrt{8}$ yielding a MT-QSL bound $T_{QSL}=\Delta\phi/\omega\simeq 0.385\pi/\omega$. The time corresponding to the brachistochrone is slightly larger, $T_{BRA}=\sqrt{11}\pi/(8\omega)\simeq 0.415\pi/\omega$. These results demonstrate that determining the path with maximum QSL efficiency is a challenging task, as the departure from the QSL typically depends on the constraints imposed on the Hamiltonian as well as on the chosen target quantum gate.
III Quantum optimal driving near the MT-quantum speed limit
To approach the quantum speed limit for unitary operators, we employ the Pontryagin Maximum Principle (PMP) [35] utilizing a cost functional that incorporates the conditions outlined in Eq. (I.1). The Pontryagin principle offers a practical optimization framework to minimize the distance of the time derivative $\dot{U}(t)$ from the plane spanned by the initial evolution operator $U_{0}=\mathbb{I}$ and the current quantum evolution operator $U(t)$.
III.1 Vaidman conditions and definition of the cost functions
We define the two following positive-valued functions:
$f_{\rm Plane}(t)$ $=$ $1-|\langle 1,U(t)\rangle|^{2}-|\langle 1,U_{\perp}(t)\rangle|^{2},$
@@ III.1 Vaidman conditions and definition of the cost functions
$f_{\rm Phase}(t)$ $=$ $\Delta H(t)^{2}{\rm Re}\left[\langle U(t),U_{0}\rangle\langle U_{0}(t),U_{\perp}(t)\rangle\right]^{2},$ (25)
@@
that quantifies the proximity of the evolution to the MT-QSL bound. The bound saturation corresponds to $f_{\rm Plane}(t)=0$ and $f_{\rm Phase}(t)=0$ at all times. The condition $f_{\rm Phase}(t)=0$ implies a phase $\beta(t)=\pm\frac{\pi}{2}$, as defined in Eq. (11). While both conditions can be simultaneously and exactly satisfied in two-dimensional quantum systems, achieving an exact solution in higher-dimensional systems depends on the available control parameters in the Hamiltonian. In such cases, a trade-off might be necessary to approximate the QSL bound by minimizing these associated quantities. Identifying the corresponding quantum path is the primary objective of the optimization algorithm described below.
@@
The goal of our quantum protocols is to achieve a target evolution operator under conditions that maximize the proximity of the Quantum Speed Limit (QSL) to its theoretical bound. To this end, we employ the optimal control formalism adapted to quantum systems [36]. It is important to note that this approach guarantees convergence to a local minimum but does not ensure that this minimum is the global one. In the following, we quantitatively assess the distance to the desired final quantum evolution operator by introducing the terminal cost function $\Phi(U(t_{f}),t_{f})=-|\langle U_{\rm target},U(t_{f})\rangle|^{2}$ which decreases when the final evolution operator gets closer to the target. Finally, to incorporate the various constraints, we define the global cost functional as follows:
$J[U(t),H(t)]$ $=$ $\Phi(U(t_{f}),t_{f})$ (26)
[…]
where the integral over $f_{0}(U(t),H(t),t)=p_{1}f_{\rm Plane}(U(t),H(t),t)+p_{2}f_{\rm Phase}(U(t),H(t),t)+p_{3}[t_{f}\Delta H(t)/\hbar]^{2}$ is the running cost. It includes two continuous components linked to the QSL conditions. The relative weights $p_{1},p_{2}$ represent a trade-off between achieving QSL saturation and obtaining a quantum operator close to the target. For $(p_{1},p_{2})=(0,0),$ the QSL optimization is completely desactivated, and the protocol is expected to deliver a final quantum gate with maximal target fidelity. We have added a small penalty term proportional to the Hamiltonian variance to suppress artificially large couplings (in all our numerical examples, we use a very small weight, $p_{3}=10^{-4}$). Identifying the set of quantum gates connected to the identity along a quantum path that closely follows the QSL (e.g., gates accessible via a protocol maintaining an average quantum speed above a certain percentage of the QSL) presents a significant challenge. Our optimization method offers a numerical approach to address this complex problem.
We next reformulate the cost function in terms of real-valued quantities, a necessary step for applying the PMP principle. To achieve this, we decompose the evolution operators and the Hamiltonian into their real and imaginary components, denoted as: $U(t)=U_{R}(t)+iU_{I}(t)$ and $H(t)=H_{R}(t)+iH_{I}(t)$. Using Eq. (2) to express the operator $U_{\perp}(t)$, we find after some basic algebra:
$f_{\rm Plane}(U(t),H(t),t)$ $=$ $1-\frac{1}{d^{2}\Delta H^{2}}\left[\left({\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{R}]\right)^{2}+\left({\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{I}]\right)^{2}\right],$
$-\frac{1}{d^{2}}\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)$
$f_{\rm Phase}(U(t),H(t),t)$ $=$ $\frac{1}{d^{4}}\left[{\rm Tr}[U_{I}]{\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]+{\rm Tr}[U_{R}]{\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)\right]^{2}.$

### v2

@@ II Saturation of the MT-QSL bound for specific families of unitary dynamics
To illustrate the constraints imposed by the Mandelstam-Tamm QSL for unitary operators, we investigate the potential saturation of this quantum speed limit in low-dimensional systems and discuss the brachistochrone approach within the QSL framework.
@@ II.1 Saturation of the MT-QSL in a 2D system
In a two-dimensional system, a generic quantum state is fully defined by two angles (up to an arbitrary phase), in agreement with the conventional Bloch sphere representation. As a result, when defining a general unitary operator, the mapping of the first state of the canonical basis corresponds to two degrees of freedom. The second basis state is necessarily mapped to an orthogonal state, but its phase can still be chosen independently. Therefore, quantum gates in a two-dimensional Hilbert space are parametrized by three degrees of freedom.
Here, we examine the saturation of the local QSL criterion (9) within a family of unitary dynamics parameterized by two control parameters $U(\theta(t),\varphi(t))=e^{-\frac{i}{2}\varphi(t)}|\psi_{+}(\theta(t))\rangle\langle 0|+e^{\frac{i}{2}\varphi(t)}|\psi_{-}(\theta(t))\rangle\langle 1|$ with the two orthogonal states $|\psi_{+}(\theta)\rangle=\cos(\theta/2)|0\rangle+\sin(\theta/2)|1\rangle$ and $|\psi_{-}(\theta)\rangle=-\sin(\theta/2)|0\rangle+\cos(\theta/2)|1\rangle$. This corresponds to a subset of the possible unitary gates in a two-dimensional system, but the extension to the more general case, which includes all possible unitary operators, is straightforward. We now determine the dynamics of the parameters $(\theta(t),\varphi(t))$ such that the QSL (9) is saturated at all times. The Hamiltonian is determined by $H(t)=i\hbar\dot{U}(t)U^{\dagger}(t)$ and is given by $H(t)=\frac{\hbar}{2}\left[\dot{\theta}U(t)\>\sigma_{y}U^{\dagger}(t)+\dot{\varphi}\>U(t)\sigma_{z}U^{\dagger}(t)\right]$ where $\sigma_{x,y,z}$ are the usual Pauli operators. We obtain $\Delta H=\frac{\hbar}{2}\sqrt{\dot{\theta^{2}}+\dot{\varphi}^{2}}.$ On the other hand, the overlap with the initial evolution operator is given by $\langle U_{0},U(t)\rangle=\cos\left(\frac{1}{2}\theta(t)\right)\cos\left(\frac{1}{2}\varphi(t)\right)$. By setting $\cos\phi=\langle U_{0},U(t)\rangle$, we find
@@
with $x=\cos\left(\frac{\theta(t)}{2}\right)$ and $y=\cos\left(\frac{\varphi(t)}{2}\right)$ (we omit the time-dependence of $x(t)$ and $y(t)$ to simplify the notation). This can be recast as $\dot{\phi}=\mathbf{a}\cdot\mathbf{b}/(2\sqrt{1-x^{2}y^{2}})$ with the column vectors $\mathbf{a}=(\dot{\theta},\dot{\varphi})^{T}$ and $\mathbf{b}=(\sqrt{1-x^{2}}y,x\sqrt{1-y^{2}})^{T}$. The Cauchy-Schwarz inequality yields
$\dot{\phi}\leq\frac{\Delta H}{\hbar}\frac{\sqrt{x^{2}+y^{2}-2x^{2}y^{2}}}{\sqrt{1-x^{2}y^{2}}}\,.$ (21)
Finally, the inequality
$\sqrt{x^{2}+y^{2}-2x^{2}y^{2}}\leq\sqrt{1-x^{2}y^{2}}$ (22)
holds because $|x|$ and $|y|$ are both within the range $[0,1]$.
The QSL is attained when both the inequalities in Eqs.  (21) and (22) are simultaneously saturated. Equation (22) saturates if either $|x|=1$ or $|y|=1$. If $|x|=1$ (meaning $\theta(t)=0$) and $0\leq|y|<1$, then the Cauchy-Schwarz inequality saturates if and only if the vectors $\mathbf{a}$ and $\mathbf{b}$ are collinear, which means $\dot{\theta}(t)=0$. Conversely, if $|y|=1$ (meaning $\varphi(t)=0$) and $0\leq x<1$, saturation occurs if and only if $\dot{\varphi}(t)=0$. To sum up, the paths saturating QSL (9) follow the great arcs of the sphere in the parameter space $(\theta,\varphi)$. In the context of quantum operators, we find a well-known result for the evolution of individual quantum states [38]: in two-dimensional systems, the geodesics associated to the Fubini-Study metric coincide with the great arcs of the Bloch sphere. To achieve QSL saturation, only one of the two angles $(\theta(t),\varphi(t))$ can evolve at a time. This example shows that saturating the QSL severely constrains the choice of possible quantum paths.
@@ II.2 Investigation of the QSL limit in 3-level systems with Rodrigues formula
In Ref. [39], a method was proposed to control the dynamics of a three-level system by separating the evolution into two components: population changes, which are parameterized using Rodrigues’ rotation formula  [Eqs.(23,24)], and phase changes. This approach was used to inversely design the Hamiltonian in a three-level system with limited access to the couplings.
Similarly, we focus here on a three-state system and the family of unitary operators $U_{r}(t)$} that govern population transfer. Specifically, we consider mappings between quantum states with positive real-valued coefficients. We denote $\mathcal{S}=\{|\psi\rangle|\forall n\in\{1,2,3\}\>\langle n|\psi\rangle\in\mathbb{R}^{+}\}$ as the corresponding set of quantum states. In other words, if $|\psi(0)\rangle=a_{1}(0)|1\rangle+a_{2}(0)|2\rangle+a_{3}(0)|3\rangle$, where $a_{n}(0)\in\mathbb{R}^{+}$ for $n\in\{1,2,3\}$, then at any time $|\psi(t)\rangle=U_{r}(t)|\psi(0)\rangle$, the condition $\langle n|\psi(t)\rangle\in\mathbb{R}^{+}$ is satisfied. Owing to their positivity, the unitary operators $U_{r}(t)$ correspond to rotations in a three-dimensional Euclidean space [39]. A 3D vector rotation is described by Rodrigues’ rotation formula:
$\mathbf{v}^{\prime}=\mathbf{v}\cos\theta+\mathbf{u}(\mathbf{u}\cdot\mathbf{v})(1-\cos\theta)+(\mathbf{u}\times\mathbf{v})\sin\theta\,$, (23)
where $\mathbf{v}$ is the vector to be rotated, $\mathbf{v}^{\prime}$ is its rotated image, $\mathbf{u}$ is the unit vector representing the axis of rotation, and $\theta$ is the angle of rotation. The corresponding relation for the quantum evolution operator is given by
@@
where we have introduced $|u\rangle=\sum_{m}u_{m}|m\rangle$ - $\mathbf{u}$ is a vector with real-valued positive coefficients and of unit norm - and $V=\sum_{i,j,k}\epsilon_{ijk}u_{i}|k\rangle\langle j|$, with $\epsilon_{ijk}$ the Levi-Civita symbol. Any evolution operator $U_{r}(t)$ that leaves invariant the subset $\mathcal{S}$, that is such that $U_{r}(t)(\mathcal{S})\subset\mathcal{S}$, can be expressed in this form.
Using this formalism, we investigate the saturation of the QSL. One of the conditions in Eq. (7) indicates that the three operators $U_{0},U(t),H(t)U(t)=i\dot{U}(t)$} must lie within a common plane. Expressing $H(t)U(t)$ as $H(t)U(t)=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U(t)$ with $(\lambda_{1}(t),\lambda_{2}(t))\in\mathbb{C}^{2}$, we derive an alternative expression for the Hamiltonian: $H(t)=\lambda_{1}(t)U^{\dagger}(t)+\lambda_{2}(t)\mathbb{I}$. Setting $a_{1,2}(t)={\rm Re}[\lambda_{1,2}(t)]$ and $b_{1,2}(t)={\rm Im}[\lambda_{1,2}(t)],$ the Hermiticity condition of the Hamiltonian implies that $a_{1}(t)(U(t)-U^{\dagger}(t))-i\>b_{1}(t)(U(t)+U^{\dagger}(t))=2ib_{2}(t)\mathbb{I}.$ Using Eq. (24) jointly with the relation $V^{\dagger}=-V,$ one finds $a_{1}(t)\sin\theta V-ib_{1}(t)\cos\theta\mathbb{I}-ib_{1}(t)(1-\cos\theta)|u\rangle\langle u|=ib_{2}(t)\mathbb{I}.$ Clearly, by linear independence of the set $\{\mathbb{I},|u\rangle\langle u|,V\}$, one has either $\theta(t)=0[2\pi]$ or $a_{1}(t)=b_{1}(t)=b_{2}(t)=0.$ In the first case, no motion occurs. In the second case the Hamiltonian $H(t)=a_{2}(t)\mathbb{I}$ merely yields a global phase factor, and the property $U_{r}(t)(\mathcal{S})\subset\mathcal{S}$ further imposes $a_{2}(t)=0$ at all times.
@@ II.3 Quantum Brachistochrone for unitary evolution and the MT-QSL limit
Indeed, this condition simply corresponds to a time-independent Schrödinger equation and is less stringent than the plane condition that arises for the MT-QSL bound, which imposes that $\dot{U}=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U(t)$.
It is important to note that a time-optimal driving in the brachistochrone framework [28] does not necessarily saturate the MT-QSL bound (9). To show this, we discuss below a few examples of time-optimal quantum gates taken from Ref. [28], and compare their duration with the minimal time from the MT-QSL bound (9).
Consider the implementation of four-state quantum gates corresponding to two interacting spins $1/2$, which dynamics follow the Hamiltonian $H(t)=-\sum_{j=x,y,z}J_{j}(t)\sigma_{j}^{1}\sigma_{j}^{2}+\sum_{a=1,2}B_{a}(t)\sigma_{z}^{a}$ with $\sigma^{1}_{j}:=\sigma_{1}\otimes\mathbf{1}_{2},$ $\sigma^{2}_{j}:=\mathbf{1}_{2}\otimes\sigma_{j}$ where $\sigma_{j}$ are the usual Pauli matrices for $j=x,y,z$, and the parameters $J_{j}(t),B_{a}(t)$ are free parameters determined by the brachistochrone equation. The time-optimal realization of a SWAP gate
@@
is reached with the time-independent Hamiltonian [28] $H_{\rm SWAP}=(\hbar\omega/\sqrt{3})[|1\rangle\langle 1|$ + $|4\rangle\langle 4|-|2\rangle\langle 2|-|3\rangle\langle 3|+2(|2\rangle\langle 3|+|3\rangle\langle 2|)]$ such that $\Delta H_{\rm SWAP}=\hbar\omega.$ As $\langle\mathbb{I},U_{\rm SWAP}\rangle=1/2,$ one has $\Delta\phi=\phi(t_{f})-\phi(0)=\pi/3$. The MT-QSL bound yields a time $T_{QSL}=\pi/(3\omega),$ which is smaller than the brachistochrone time [28] $T_{BRA}=\sqrt{3}\pi/(4\omega)\simeq 0.433\pi/\omega$. Similarly, when considering the “QFT” gate (related to the quantum Fourier transform gate $U_{\rm QFT}$ with Hadamard gates (see [40])),
we obtain $|\langle\mathbb{I},U_{\rm QFT}\rangle|=1/\sqrt{8}$ yielding a MT-QSL bound $T_{QSL}=\Delta\phi/\omega\simeq 0.385\pi/\omega$. The time corresponding to the brachistochrone [28] is slightly larger, $T_{BRA}=\sqrt{11}\pi/(8\omega)\simeq 0.415\pi/\omega$. These results demonstrate that determining the path with maximum QSL efficiency is a challenging task, as the departure from the QSL typically depends on the constraints imposed on the Hamiltonian as well as on the chosen target quantum gate.
III Quantum optimal driving near the MT-quantum speed limit: geometric optimization with Pontryagin Maximum’s principle
To approach the quantum speed limit for unitary operators, we employ the Pontryagin Maximum Principle (PMP) [41] using a cost functional that incorporates the conditions outlined in Eq. (I.1). The Pontryagin principle offers a practical optimization framework to minimize the distance of the time derivative $\dot{U}(t)$ from the plane spanned by the initial evolution operator $U_{0}=\mathbb{I}$ and the current quantum evolution operator $U(t)$. We consider quantum protocols of fixed total duration $t_{f}$.
III.1 Vaidman conditions and definition of the cost functionals
@@
$|\dot{\phi}(t)|=\frac{\Delta H}{\hbar}\sqrt{1-f_{\rm Plane}(U,H)}\sqrt{1-f_{\rm Phase}(U,H)}.$ (29)
For lighter notation, we have omitted the explicit time dependence on the right-hand side. These functions can be expressed as:
$f_{\rm Plane}(U(t),H(t))$ $=$ $\frac{1-|\langle\mathbb{I},U(t)\rangle|^{2}-|\langle\mathbb{I},U_{\perp}(t)\rangle|^{2}}{1-|\langle\mathbb{I},U(t)\rangle|^{2}},$ (30)
@@
$f_{\rm Phase}(U(t),H(t))$ $=$ $\frac{{\rm Re}\left[\langle U(t),\mathbb{I}\rangle\langle\mathbb{I},U_{\perp}(t)\rangle\right]^{2}}{|\langle U(t),\mathbb{I}\rangle|^{2}|\langle\mathbb{I},U_{\perp}(t)\rangle|^{2}},$ (31)
@@
with $U_{\perp}(t)$ defined from Eq.(2). The expression holds as long as $|\langle\mathbb{I},U(t)\rangle|\neq 1$ and $|\langle U(t),\mathbb{I}\rangle||\langle\mathbb{I},U_{\perp}(t)\rangle|\neq 0$. The saturation of the QSL bound corresponds to $f_{\rm Plane}(U(t),H(t))=0$ and $f_{\rm Phase}(U(t),H(t))=0$ at all times. While both conditions can be simultaneously and exactly satisfied at all times in two-dimensional quantum systems, achieving an exact solution in higher-dimensional systems depends on the available control parameters in the Hamiltonian. In such cases, a trade-off might be necessary to approximate the QSL bound by minimizing both $f_{\rm Plane}$ and $f_{\rm Phase}$. Identifying the corresponding quantum path is the primary objective of the optimization algorithm described below.
@@
The goal of our quantum protocols is to achieve a target evolution operator under conditions that maximize the proximity of the Quantum Speed Limit (QSL) to its theoretical bound. To this end, we employ the optimal control formalism adapted to quantum systems [42]. It is important to note that this approach guarantees convergence to a local minimum but does not ensure that this minimum is the global one. In the following, we quantitatively assess the distance to the desired final quantum evolution operator by introducing the infidelity as terminal cost function $\Phi(U(t_{f}),t_{f})=1-|\langle U_{\rm target},U(t_{f})\rangle|^{2}$, which decreases as the final evolution operator gets closer to the target. Finally, to incorporate the various constraints, we define the global cost functional as follows:
$J[U,H]$ $=$ $\Phi(U(t_{f}),t_{f})$ (32)
$+$ $\frac{1}{t_{f}}\int_{0}^{t_{f}}dt^{\prime}f_{0}(U(t^{\prime}),H(t^{\prime}),t^{\prime})$
The integral over the running cost $f_{0}(U(t),H(t),t)=p_{1}f_{\rm Plane}(U(t),H(t),t)+p_{2}f_{\rm Phase}(U(t),H(t),t)$ yields the QSL contribution to the cost functional $J[U,H]$. Detailed expressions for the running cost functions $f_{\rm Plane}(U(t),H(t),t)$ and $f_{\rm Phase}(U(t),H(t),t)$, in terms of the real and imaginary parts of the evolution operator and Hamiltonian, are provided in Appendix A.
The cost functional $J[U,H]$ includes two continuous components linked to the QSL conditions, namely $J_{\rm plane}[U,H]=\frac{1}{t_{f}}\int_{0}^{t_{f}}f_{\rm plane}(U(t),H(t))dt$ and $J_{\rm phase}[U,H]=\frac{1}{t_{f}}\int_{0}^{t_{f}}f_{\rm phase}(U(t),H(t))dt$. Their respective weights $p_{1}$ and $p_{2}$ represent a trade-off between achieving QSL saturation and obtaining a quantum operator close to the target. For $(p_{1},p_{2})=(0,0),$ the QSL optimization is completely deactivated, and the protocol is expected to deliver a final quantum gate with maximal target fidelity. The plane and phase running cost functions $f_{\rm Plane}(U(t),H(t))$ and $f_{\rm Phase}(U(t),H(t))$ play a symmetric role in Eq. (29). A small mismatch in either the plane or phase condition ($f_{\rm Plane}(U,H)=\delta$ or $f_{\rm Phase}(U,H)=\delta$) results in an identical degradation of the instantaneous quantum speed. This suggests that optimal results are obtained when equal weights ($p_{1}=p_{2}$) are attributed to the two cost functionals $J_{\rm Plane}[U,H]$ and $J_{\rm Phase}[U,H]$.
Identifying the set of quantum gates connected to the identity along a quantum path that closely follows the QSL (e.g., gates accessible via a protocol maintaining an average quantum speed above a certain percentage of the QSL) presents a significant challenge. Our optimization method offers a numerical approach to address this complex problem.